Workers: Akutsu, Simon, Miyo
- Assembled additional two photosensor heads; now we have all we required; in other words, three in total.
- Continued to make a test bench to calibrate the sensor heads.
- Connected a sensor head to the burndy connector P23 in the chamber with 3-m long patch cables.
Seemed noisy so far... Details will be reported later
- At first we used an OSEM sat box to drive the photosensor, but the reflected light amount seemed low, so tried to increase the emitting power from the LED by changing the power supply to the LED from the OSEM sat box's one to just a power supply. According to the spec sheet the max rating is around 100mA, and so we like to use 70mA. But still noisy...
Issue: how should we connect the coil pins and the electric wires...?
[Ongoing] Need cares of water springs from anchor bolts for oplev pylons.
[Ongoing] Need to confirm what are the scratch lines made by the measurement by the KEK team last week.
- Calibrate the PS.
- Complete the calibration of the oplev; the arm length was reported by Simon, so....
- Check the balance of the suspended WAB with some wires and sensors are attached.
- Check how the damping would be now after the replacement of the damping magnets.
- Check the reproducibility of the alignment of the suspension among several modes.
- Prepare calibration of photosensors.
- Check heights of several points.
- Take care of water spring??
- Two burndy connectors, vacuum compatible.
Stages outside the chamber to ease wiping our legs.
Good rail(?) to move the WAB suspension in the cryostat in the main beam direction.
Chain block? to support the load in the chamber.
- A clock to know the time easily.
- M6 screw set.